OpenStudy (anonymous):

Find the slope of the tangent line to the curve y=logBase5(3x-1) at point x=3

OpenStudy (raden):

it can rewriten as 5^y = 3x-1 take derivative by implicit 5^y ln5 dy = 3 dx dy/dx = 3/(5^y ln 5) so, m = 3/(5^y ln 5) = 3/((3x-1)ln5) now, put the value of x=3 m = 3/(3(3)-1)ln5 = 3/8ln5

OpenStudy (anonymous):

Thank you!(:

OpenStudy (raden):

you'r welcome